Each of two brothers can wash a car in 1 hour; however their sister can wash the car in 45 minutes.If all three work together, how long will it take to wash the car?
Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(7384)

Each of two brothers can wash a car in 1 hour; however their sister can wash the car in 45 minutes.If all three work together, how long will it take to wash the car?
.
Rate of each brother: 1 car per 1 hour = 1/1
Rate of sister: 1 car per 45 min = 1 car per .75 hour = 1/.75
.
Let x = time (hours) for all three working together
then
x(1/1 + 1/1 + 1/.75) = 1
x(1 + 1 + 1/.75) = 1
multiplying both sides by .75:
x(.75 + .75 + 1) = .75
x(2.5) = .75
x = .75/2.5
x = .3 hour
or
x = 18 minutes

Answer by stanbon(75887)

Each of two brothers can wash a car in 1 hour; however their sister can wash the car in 45 minutes.
If all three work together, how long will it take to wash the car?
---
Equation:
rate + rate + rate = together rate
1 + 1 + (1/(3/4)) = 1/x
-----
2 (4/3) = 1/x
---
10/3 = 1/x
x = 3/10 hr
---
(3/10)hr = (3/10)60 min = 18 minutes
---
They can do it together in 18 minutes.
Cheers,
Stan H.